The Chunnel connects the UK and Fance by a railway tunnel under the English Channel. The British started tunneling 2.5 months before the French and averaged 0.63 kilometer per month. The French averaged 0.47 kilometer per month. When the two sides met, they had tunneled 37.9 kilometers.
What is the length of tunnel each country built?

Answer by stanbon(48558)

The Chunnel connects the UK and Fance by a railway tunnel under the English Channel. The British started tunneling 2.5 months before the French and averaged 0.63 kilometer per month. The French averaged 0.47 kilometer per month. When the two sides met, they had tunneled 37.9 kilometers.
-----------------
Distance worked together = 37.9 - 2.5*0.63 = 36.325 km
EQUATION:
0.63x + 0.47x = 36.325 km
1.1x = 36.325 km
x = 33.0227... months
=====================
Cheers,
Stan H.